OpenStudy (anonymous):

2x - 9/10 <= -7 or 2x - 9/10 >= 4 What is the solution of the compound inequality?

OpenStudy (saifoo.khan):

did u gave it a try?

OpenStudy (anonymous):

I did...but I am just horrible at fractions. I am lost.

OpenStudy (anonymous):

No way I will get this one. I have a headache already. LOL

OpenStudy (anonymous):

I know it is {x|x <= ?? OR x >= ??}

OpenStudy (saifoo.khan):

alright, let me solve!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

ew nasty quadratic.

OpenStudy (saifoo.khan):

umm, no not quadratic my friend.

OpenStudy (anonymous):

>.> how are you solving this?

OpenStudy (saifoo.khan):

this will be solved separately. 2x - 9/10 <= -7 & 2x - 9/10 >= 4

OpenStudy (saifoo.khan):

\[\Huge {x|x \le -\frac{61}{11} or x \ge \frac{23}{9} }\]

OpenStudy (anonymous):

what is that notation called again?

OpenStudy (saifoo.khan):

interval notation.

OpenStudy (anonymous):

ah yes.

OpenStudy (anonymous):

How come I got -61/20?

OpenStudy (anonymous):

not correct. Its says to carefully check over the work tht you did to solve each inequality.

OpenStudy (saifoo.khan):

oh, thts a typo... mybrains is right.

OpenStudy (anonymous):

;)

OpenStudy (saifoo.khan):

i hate when my calc does that! :@

OpenStudy (saifoo.khan):

whts your name mybrain?

OpenStudy (anonymous):

Alan

OpenStudy (saifoo.khan):

Al-an?

OpenStudy (anonymous):

my nickname refers to my constant struggle with my brain haha

OpenStudy (saifoo.khan):

lol, jk.

OpenStudy (saifoo.khan):

haha!

OpenStudy (anonymous):

I tried -61/20 or 23/9 and it is not correct.

OpenStudy (anonymous):

how are you checking it dumbbunny?

OpenStudy (saifoo.khan):

dang it.

OpenStudy (saifoo.khan):

the second one is: 49/20

OpenStudy (anonymous):

Well, I am not. Fractions give me fits. I just put in the two answers you guys came up with and it says "Sorry that is not correct."

OpenStudy (saifoo.khan):

still? >.< :(

OpenStudy (anonymous):

so I got -61/20 or 49/20. Those are the solutions but you should also be saying x <= -61/20 or x>= 49/20

OpenStudy (anonymous):

You don't need to be good with fractions. Believe me I still can't add them on paper, just put them in your calculator.

OpenStudy (saifoo.khan):

hmm. right^

OpenStudy (anonymous):

LOL, well you were both close...the correct answer is -61/2 and 49/2 but you both tried very hard and I appreciate that.

OpenStudy (anonymous):

LOL what. I feel cheated, how could it be just 2?? did you write the probelm down wrong :P?

OpenStudy (anonymous):

No, I triple checked it.

OpenStudy (saifoo.khan):

i think so..

OpenStudy (anonymous):

I gave you a medal anyway.

OpenStudy (anonymous):

:]

OpenStudy (saifoo.khan):

hehe, thnks!